OpenStudy (anonymous):

One serving of water is considered 8 fl oz. ,It is recommended that adults drink 8 servings of water per day. How many quarts of water is this?

OpenStudy (johnweldon1993):

Taking into mind the fact that 1 serving = 8 fl oz so 8 servings = 8 * 8 fl oz 8 servings = 64 fl oz Now...it asks for quarts...not fluid ounces. remember that 1 US fluid ounce = 0.03125 US quarts so 64 fluid ounces = 64 * 0.03125 quarts What does that equal?